Suppose P(A) = 0.25. The probability of complement of A is:

A. 0.82
B. 0.50
C. 0.75
D. 0.25
Answer #1

Solution: Since we know that

P(A)+P(A^{c})=1; where A^{c} is a complement of A.

Now, given that

P(A) =0.25.

Therefore,

0.25+P(A^{c})=1

P(A^{c})=1-0.25=0.75

Hence,

(c) 0.75 is correct